OpenStudy (anonymous):

Calculate the Fermi energy at 300 K for germanium containing 5x1022 arsenic atoms /m3. Let Eg (Ge) = 0.66 eV

OpenStudy (aliqanber):

\[E _{f} = \frac{ \hbar^{2} }{ 2m } \left( \frac{ 3 \pi^{2} N }{ V } \right)^{2/3}\]

OpenStudy (dboome):

\[E_F(T) = E_F(0)(1-\frac{ \pi^2k^2T^2 }{ 12E_F(0) } )\]The fermi energy at temperature T can be expressed in terms of the fermi energy at 0K, E_F(0). which is the equation aliqanber gave you. Remember that you already have N/V, which is the density you are given.